# High-dimensional sliced inverse regression with endogeneity

### v1

@@ Acknowledgments
$\leq\left\|\hat{\bm{\eta}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}\right\|_{\infty}+\left\|\left(\bm{\Sigma}_{\widehat{\mathbf{X}}}-\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\right)\tilde{\bm{\beta}}\right\|_{\infty}.$ (23)
@@
We will upper bound each term on the right-hand side of (23). For the first term, note by definition, $\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}=\tilde{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\widehat{\bm{\eta}}$ i.e., the eigenvector associated with the largest eigenvalue of $\bm{\Lambda}=\operatorname{\text{Var}}\left\{E(\widehat{\mathbf{X}}\mid y)\right\}$. Hence, we obtain
$\left\|\widehat{\bm{\eta}}-\bm{\Sigma}_{{\widehat{\mathbf{X}}}}\tilde{\bm{\beta}}\right\|_{\infty}=\left\|\widehat{\bm{\eta}}-\tilde{\bm{\eta}}\right\|_{\infty}$ $\leq\left\lVert\widehat{\bm{\eta}}-\tilde{\bm{\eta}}\right\rVert_{2}$
=$\stackrel{{\scriptstyle(i)}}{{=}}\sqrt{1-\hat{\rho}^{2}}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}C\sqrt{\frac{L}{\lambda}\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}},$
for a large positive constant $C$, where step (i) above follows since
$\left\lVert\widehat{\bm{\eta}}-\tilde{\bm{\eta}}\right\rVert_{2}^{2}=(\widehat{\bm{\eta}}-\tilde{\bm{\eta}})^{\top}(\widehat{\bm{\eta}}-\tilde{\bm{\eta}})$ $=\widehat{\bm{\eta}}^{\top}\widehat{\bm{\eta}}-2\widehat{\bm{\eta}}^{\top}\tilde{\bm{\eta}}+\tilde{\bm{\eta}}^{\top}\tilde{\bm{\eta}}$
$=1-2(\widehat{\bm{\eta}}^{\top}\bm{\eta})(\bm{\eta}^{\top}\widehat{\bm{\eta}})+\widehat{\bm{\eta}}^{\top}\bm{\eta}\bm{\eta}^{\top}\bm{\eta}\bm{\eta}^{\top}\widehat{\bm{\eta}}$
$=1-2\hat{\rho}^{2}+\hat{\rho}^{2}=1-\hat{\rho}^{2},$
and step (ii) follows from applying Lemma 2.
@@
For the second term on the right-hand side of (23), using (42) we have
@@
Hence with the choice of $\mu_{2}$ as stated in the theorem, we have
@@
$\left\|P(\hat{\bm{\beta}})-P({\bm{\beta}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\sqrt{\dfrac{L_{1}}{\lambda}}+L_{1}\right),~{}L_{1}=L\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}$
@@
$\dfrac{1}{n}\hat{\mathcal{X}}^{\top}_{\hat{S}}(\tilde{\mathbf{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}})=\mu_{2}\text{sgn}(\hat{\bm{\beta}}_{\hat{S}}),$ (25)
@@
$\mu_{2}\text{sgn}(\hat{\bm{\beta}}_{\hat{S}})=\dfrac{1}{n}\hat{\mathcal{X}}^{\top}_{\hat{S}}(\tilde{\mathbf{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}})$ $=\hat{\bm{\eta}}_{S}-\dfrac{1}{n}\hat{\mathcal{X}}_{S}^{\top}\hat{\mathcal{X}}_{S}\hat{\bm{\beta}}_{S}$
@@
$\hat{\bm{\beta}}_{S}-\tilde{\bm{\beta}}_{S}=\widehat{\mathbf{C}}_{SS}^{-1}\left\{(\hat{\bm{\eta}}_{S}-\widehat{\mathbf{C}}_{SS}\tilde{\bm{\beta}}_{S})-\mu_{2}\text{sgn}(\hat{\bm{\beta}}_{\hat{S}})\right\},$ (30)
@@
and by the choice of $\mu_{2}$ as stated in the theorem. Thus we have $\text{sgn}(\hat{\bm{\beta}}_{S})=\text{sgn}(\tilde{\bm{\beta}}_{S})$, and since $\hat{\bm{\beta}}_{S^{c}}=\bm{0}$ by construction then we have $\hat{S}=S$.
@@
$\dfrac{1}{n}\hat{\mathcal{X}}^{\top}_{\hat{S}^{c}}(\tilde{\mathbf{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}})$ $=\hat{\bm{\eta}}_{S^{c}}-\dfrac{1}{n}\hat{\mathcal{X}}_{S^{c}}^{\top}\hat{\mathcal{X}}_{S}\hat{\bm{\beta}}_{S}$
@@
$\left\|\dfrac{1}{n}\hat{\mathcal{X}}^{\top}_{\hat{S}^{c}}(\tilde{\mathbf{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}})\right\|_{\infty}$ $\leq\left\|\hat{\bm{\eta}}_{S^{c}}-\dfrac{1}{n}\hat{\mathcal{X}}_{S^{c}}^{\top}\hat{\mathcal{X}}_{S}\tilde{\bm{\beta}}_{S}\right\|_{\infty}+\|\widehat{\mathbf{C}}_{S^{c}S}\widehat{\mathbf{C}}_{SS}^{-1}\|_{\infty}\|(\hat{\bm{\eta}}_{S}-\widehat{\mathbf{C}}_{SS}\tilde{\bm{\beta}}_{S})-\mu_{2}\text{sgn}(\hat{\bm{\beta}}_{{S}})\|_{\infty}$
@@
S2.9 Proof of Theorem 4
Consider the multiple index model with $\mathbf{B}\in\mathbb{R}^{p\times d}$. In this case, the quantities $\bm{\eta}$ and $\widehat{\bm{\eta}}$ are the $p\times d$ semi-orthogonal matrices of eigenvectors of $\bm{\Lambda}$ and $\widehat{\bm{\Lambda}}$, respectively. Let $\widetilde{\bm{\eta}}=P(\bm{\eta})\widehat{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\widehat{\bm{\eta}}$ be the projection of $\widehat{\bm{\eta}}$ on $\bm{\eta}$. By Lemma 1, we have $\Sigma_{\widehat{\mathbf{X}}}\mathbf{B}=\bm{\eta}\widetilde{\mathbf{D}}$ where $\widetilde{\mathbf{D}}$ is a non-zero $d\times d$ diagonal matrix, such that $\bm{\eta}=\Sigma_{\widehat{\mathbf{X}}}\mathbf{B}\widetilde{\mathbf{D}}^{-1}$. Hence, we can write
@@
where $\tilde{\bm{\beta}}=\widetilde{\mathbf{D}}^{-1}\bm{\eta}^{\top}{\widehat{\bm{\eta}}}$. When $\mathbf{H}=\bm{\eta}^{\top}{\widehat{\bm{\eta}}}$ is a $d\times d$ invertible matrix, the column space spanned by $\tilde{\mathbf{B}}$ and that spanned by $\mathbf{B}$ are the same, and so are the index sets for their non-zero rows.
@@
$\left\lVert\hat{\bm{\beta}}_{k}-\tilde{\bm{\beta}}_{k}\right\rVert_{2}\leq\dfrac{3}{\kappa_{2}}\mu_{2k}\sqrt{s},$
for any $\mu_{2k}\geq 2\left\lVert\bm{\eta}_{k}-n^{-1}\widehat{\mathcal{X}}^{\top}\widehat{\mathcal{X}}\widetilde{\bm{\beta}}_{k}\right\rVert_{\infty}$. Hence, by a similar argument to the proof of Theorem 2, and with the choice of $\mu_{2}$ as stated in the theorem, there exist constants $C,C^{\prime}$, and $C^{\prime\prime}$ such that
$\left\|P(\hat{\bm{\beta}})-P({\bm{\beta}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\sqrt{\dfrac{L_{1}}{\lambda_{d}}}+L_{1}\right),~{}L_{1}=L\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}$
Finally, turning to variable selection consistency, let $\mathcal{S}_{k}=\left\{k:\beta_{jk}\neq 0\right\}$ and $\widehat{\mathcal{S}}_{k}=\left\{k:\hat{\beta}_{jk}\neq 0\right\}$. By a similar argument to that in the proof of Theorem 3, under the condition on $\mu_{k}$ as stated in Theorem 4, we have $\widehat{\mathcal{S}}_{k}=\mathcal{S}_{k}$ for $k=1,\ldots,d$ with probability at least $1-\exp(C^{\prime}\log q)-\exp(C^{\prime\prime}\log pq)$. Note the diagonal element of the estimated projection matrix $[\mathcal{P}(\widehat{\mathbf{B}})]_{jj}=0$ if and only if all one $\hat{\beta}_{jk}=0$ across $k=1,\ldots,d$. Hence, taking the union bound, we have $\mathcal{S}=\mathcal{S}$ with a probability at least $1-d\exp(C^{\prime}\log q)-d\exp(C^{\prime\prime}\log pq)$.
S3 Additional derivations
S3.1 Proof of Lemma 4

### v2

@@ B.6 Proof of Theorem 2
$\leq\underbrace{\left\|\hat{\bm{\eta}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}\right\|_{\infty}}_{\text{I}}+\underbrace{\left\|\left(\bm{\Sigma}_{\widehat{\mathbf{X}}}-\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}\right)\tilde{\bm{\beta}}\right\|_{\infty}}_{\text{II}}.$ (25)
We will upper bound each term on the right-hand side of (25).
@@
Bounding (I): Recall that by definition $\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}=\tilde{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\widehat{\bm{\eta}}$ i.e., where $\bm{\eta}$ is the eigenvector associated with the largest eigenvalue of $\bm{\Lambda}=\operatorname{\text{Var}}\left\{E(\widehat{\mathbf{X}}\mid y)\right\}=\bm{\Gamma}^{\top}\bm{\Omega}\bm{\Gamma}$ with $\bm{\Omega}=\operatorname{\text{cov}}\left\{E(\mathbf{Z}\mid y)\right\}$. Hence, we have $\|\hat{\bm{\eta}}-\bm{\Sigma}_{\widehat{\mathbf{X}}}\tilde{\bm{\beta}}\|_{\infty}=\|\widehat{\bm{\eta}}-\tilde{\bm{\eta}}\|_{\infty}$.
Following the same notation as in Lin et al., (2019), let $\mathcal{Z}_{H}$ be the $H\times q$ matrix where the $h$th row corresponds to the sample mean of $\mathbf{Z}$ within the $h$th slice, for $h=1,\ldots,H$, and let $\widehat{\mathcal{X}}_{H}=\mathcal{Z}_{H}\widehat{\bm{\Gamma}}$ be the corresponding matrix of the sliced mean for $\widehat{\mathbf{X}}$. Hence, we can project $\widehat{\mathcal{X}}_{H}$ into the column space of $\bm{\Lambda}$, i.e $\text{span}(\bm{\Gamma}^{\top}\bm{\eta}(\bm{\Omega}))$, and decompose
$\widehat{\mathcal{X}}_{H}=\mathcal{V}_{H}+\mathcal{T}_{H},$
where $\mathcal{V}_{H}=\widehat{\mathcal{X}}_{H}\mathbf{P}$ and $\mathcal{T}_{H}=\widehat{\mathcal{X}}_{H}(\mathbf{I}_{p}-\mathbf{P})$, and $\mathbf{P}=\bm{\Gamma}^{\top}\bm{\eta}(\bm{\Omega})\{\bm{\eta}(\bm{\Omega})^{\top}\bm{\Gamma}\bm{\Gamma}^{\top}\bm{\eta}(\bm{\Omega})\}^{-1}\bm{\eta}(\bm{\Omega})^{\top}\bm{\Gamma}$ is the $p\times p$ projection matrix into the subspace spanned by $\bm{\Gamma}^{\top}\bm{\eta}(\bm{\Omega})$. By the same argument as in Lin et al., (2019), we have
$\|\hat{\bm{\eta}}-\tilde{\bm{\eta}}\|_{\infty}\leq\dfrac{1}{\sqrt{\hat{\lambda}}}\|\mathcal{T}_{H}\|_{\max},$ (26)
where $\hat{\lambda}$ is the largest eigenvalue for $\widehat{\bm{\Lambda}}=H^{-1}\widehat{\mathcal{X}}_{H}^{\top}\widehat{\mathcal{X}}_{H}=H^{-1}\widehat{\bm{\Gamma}}^{\top}\mathcal{Z}_{H}^{\top}\mathcal{Z}_{H}\widehat{\bm{\Gamma}}$. We next bound $\|\mathcal{T}_{H}\|_{\max}$. To do it, we also project $\mathcal{Z}_{H}$ into the column space of $\bm{\Omega}$, and write
$\mathcal{Z}_{H}=\mathcal{W}_{H}+\mathcal{U}_{H},$ (27)
where $\mathcal{W}_{H}$ and $\mathcal{U}_{H}$ are similarly, the projection of $\mathcal{Z}_{H}$ to $\text{span}(\bm{\eta}(\bm{\Omega}))$ and its orthogonal complement. Applying the results from Lin et al., (2019), we have $\|\mathcal{U}_{H}\|_{\max}=O_{p}\left(\sqrt{\log(q)/n}\right)$. With this result, we obtain
$\mathcal{T}_{H}=\widehat{\mathcal{X}}_{H}(\mathbf{I}_{p}-\mathbf{P})$ $=\mathcal{Z}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})=\mathcal{W}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})+\mathcal{U}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})$ (28)
@@
We consider each term on the right-hand side of (28) separately. For the first term, we have
$\mathcal{W}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})=\mathcal{W}_{H}(\widehat{\bm{\Gamma}}-{\bm{\Gamma}})(\mathbf{I}_{p}-\mathbf{P})+\mathcal{W}_{H}{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\stackrel{{\scriptstyle(i)}}{{=}}\mathcal{W}_{H}(\widehat{\bm{\Gamma}}-{\bm{\Gamma}})(\mathbf{I}_{p}-\mathbf{P}),$ (29)
where step $(i)$ follows from the fact that $\mathcal{W}_{H}\in\text{span}(\bm{\eta}(\bm{\Omega}))$ and hence $\mathcal{W}_{H}{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})=\bm{0}$. Therefore, we have
$\|\mathcal{W}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\|_{\max}=\|\mathcal{W}_{H}(\widehat{\bm{\Gamma}}-{\bm{\Gamma}})(\mathbf{I}_{p}-\mathbf{P})\|_{\max}$ $\leq\|\mathcal{W}_{H}\|_{\max}\|(\widehat{\bm{\Gamma}}-{\bm{\Gamma}})(\mathbf{I}_{p}-\mathbf{P})\|_{1}$
$\leq\|\mathcal{W}_{H}\|_{\max}\|\widehat{\bm{\Gamma}}-\bm{\Gamma}\|_{1}\|\mathbf{I}_{p}-\mathbf{P}\|_{1},$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}C\sqrt{\log(q)}\|\mathcal{W}_{H}\|_{\max}\|\widehat{\bm{\Gamma}}-\bm{\Gamma}\|_{1},$
where step $(ii)$ follows from the triangle inequality $0<\|\mathbf{I}_{p}-\mathbf{P}\|_{1}\leq\|\mathbf{I}_{p}\|_{1}+\|\mathbf{P}\|_{1}=1+\Psi_{1}$. Furthermore,
$\|\mathcal{W}\|_{\max}=\|\mathcal{Z}_{H}\bm{\eta}(\bm{\Omega})\bm{\eta}(\bm{\Omega})^{\top}\|_{\max}\leq\|\mathcal{Z}_{H}\|_{\max}\|\bm{\eta}(\bm{\Omega})\bm{\eta}(\bm{\Omega})^{\top}\|_{1}=O_{p}(\Psi_{2}\sqrt{\log q}).$
Hence, we have $\|\mathcal{W}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\|_{\max}\leq C\Psi_{2}(1+\Psi_{1})\|\widehat{\bm{\Gamma}}-\bm{\Gamma}\|_{1}$. Using the result from Theorem 3.1, we obtain
$\|\mathcal{W}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\|_{\max}\leq C\Psi_{2}(1+\Psi_{1})r\sqrt{\frac{\log pq}{n}}.$ (30)
Now, for the second term in (23), by a similar argument we have
$\|\mathcal{U}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\|_{\max}\leq\|\mathcal{U}_{H}\|_{\max}\|\widehat{\bm{\Gamma}}\|_{1}\leq C(1+\Psi_{1})\sqrt{\dfrac{\log(q)}{n}}\left(\|\widehat{\bm{\Gamma}}-\bm{\Gamma}\|_{1}+\|\bm{\Gamma}\|_{1}\right),$ (31)
where the last step follows from the triangle inequality. Since $\|\bm{\Gamma}\|_{1}\leq L$, we have
$\|\mathcal{U}_{H}\widehat{\bm{\Gamma}}(\mathbf{I}_{p}-\mathbf{P})\|_{\max}\leq C(1+\Psi_{1})\left(L\sqrt{\dfrac{\log q}{n}}+r\sqrt{\frac{\log q(\log pq)}{n^{2}}}\right).$ (32)
Substituting (30) and (32) into (28) and applying (26), we obtain
[…]
using the assumption that $\log(q)\log(pq)=o(n^{2})$. By proposition 3, there exists a positive constant $C$ such that $\lambda(\bm{\Omega})/\hat{\lambda}(\bm{\Omega})\leq C$. Combining it with the inequality (20), we obtain
$\dfrac{\lambda}{\hat{\lambda}}\leq\dfrac{\lambda(\bm{\Omega})}{\hat{\lambda}(\bm{\Omega})}\times\dfrac{\nu_{\max}(\bm{\Gamma})}{\nu_{\min}(\widehat{\bm{\Gamma}})}\leq C,$
with a probability of at least $1-\exp\left(-C^{\prime}\log q\right)-\exp(-C^{\prime\prime}\log pq)$, where the last inequality follows from condition 4.2 and with probability one, the matrix $\widehat{\bm{\Gamma}}$ obtained from the first stage has full rank. Hence, with the same probability, we have
$\|\hat{\bm{\eta}}-\tilde{\bm{\eta}}\|_{\infty}\leq\dfrac{1}{\sqrt{\hat{\lambda}}}\|\mathcal{T}\|_{\max}\leq C(1+\Psi_{1})\left(L\sqrt{\dfrac{\log q}{n\lambda}}+\Psi_{2}r\sqrt{\dfrac{\log pq}{n\lambda}}\right).$
Bounding (II): We now bound the second term of (25). Using (52) in the proof of Proposition 3, we obtain
@@
Finally, with the choice of $\mu_{2}$ as stated in the theorem, we have
@@
$\left\|P(\hat{\bm{\beta}})-P({\bm{\beta}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right),$
where
$\rho_{1}=L^{2}+\dfrac{L(1+\Psi_{1})}{\sqrt{\lambda}},~\rho_{2}=L\sigma_{\max}+\dfrac{(1+\Psi_{1})\Psi_{2}}{\sqrt{\lambda}}$
@@ B.7 Proof of Theorem 3
[…]
@@
[…]
@@
[…]
@@
and by the choice of $\mu_{2}$ as stated in the theorem. Thus we have $\text{sgn}(\hat{\bm{\beta}}_{S})=\text{sgn}(\tilde{\bm{\beta}}_{S})$, and since $\hat{\bm{\beta}}_{S^{c}}=0$ by construction then we have $\hat{S}=S$.
@@
[…]
@@
[…]
@@
B.8 Proof of Theorem 4
Consider the multiple index model with $\mathbf{B}\in\mathbb{R}^{p\times d}$. In this case, the quantities $\bm{\eta}$ and $\widehat{\bm{\eta}}$ are the $p\times d$ semi-orthogonal matrices of eigenvectors of $\bm{\Lambda}$ and $\widehat{\bm{\Lambda}}$, respectively. Let $\widetilde{\bm{\eta}}=P(\bm{\eta})\widehat{\bm{\eta}}=\bm{\eta}\bm{\eta}^{\top}\widehat{\bm{\eta}}$ be the projection of $\widehat{\bm{\eta}}$ on $\bm{\eta}$. By Lemma 1, we have $\bm{\Sigma}_{\widehat{\mathbf{X}}}\mathbf{B}=\bm{\eta}\widetilde{\mathbf{D}}$ where $\widetilde{\mathbf{D}}$ is a non-zero $d\times d$ diagonal matrix, such that $\bm{\eta}=\Sigma_{\widehat{\mathbf{X}}}\mathbf{B}\widetilde{\mathbf{D}}^{-1}$. Hence, we can write
@@ B.8 Proof of Theorem 4
where $\widetilde{\mathbf{B}}=\widetilde{\mathbf{D}}^{-1}\bm{\eta}^{\top}{\widehat{\bm{\eta}}}$. When $\mathbf{H}=\bm{\eta}^{\top}{\widehat{\bm{\eta}}}$ is a $d\times d$ invertible matrix, the column space spanned by $\tilde{\mathbf{B}}$ and that spanned by $\mathbf{B}$ are the same, and so are the index sets for their non-zero rows.
@@
$\lVert\hat{\bm{\beta}}_{k}-\tilde{\bm{\beta}}_{k}\rVert_{2}\leq\dfrac{3}{\kappa_{2}}\mu_{2k}\sqrt{s},$
for any $\mu_{2k}\geq 2\lVert\bm{\eta}_{k}-n^{-1}\widehat{\mathcal{X}}^{\top}\widehat{\mathcal{X}}\widetilde{\bm{\beta}}_{k}\rVert_{\infty}$. Hence, by a similar argument to the proof of Theorem 2, and with the choice of $\mu_{2}$ as stated in the theorem, there exist constants $C,C^{\prime}$, and $C^{\prime\prime}$ such that
$\left\|P(\hat{\bm{\beta}}_{k})-P({\bm{\beta}_{k}})\right\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right),$
where
$\rho_{1}=L^{2}+\dfrac{L(1+\Psi_{1})}{\sqrt{\lambda_{d}}},~\rho_{2}=L\sigma_{\max}+\dfrac{(1+\Psi_{1})\Psi_{2}}{\sqrt{\lambda}_{d}}$
with probability at least $1-\exp(C^{\prime}\log q)-\exp(C^{\prime\prime}\log pq)$. Hence, by the same argument as in the proof of Theorem 2 in Lin et al., (2019), we also have
$\|P(\widehat{\mathbf{B}})-P(\mathbf{B})\|_{F}\leq C\frac{\sqrt{s}}{\kappa_{2}}\left(\rho_{1}\sqrt{\frac{\log q}{n}}+\rho_{2}r\sqrt{\frac{\log pq}{n}}\right),$
with probability at least $1-d\exp(C^{\prime}\log q)-d\exp(C^{\prime\prime}\log pq)$ as claimed.
Finally, turning to variable selection consistency, let $\mathcal{S}_{k}=\left\{k:\beta_{jk}\neq 0\right\}$ and $\widehat{\mathcal{S}}_{k}=\left\{k:\hat{\beta}_{jk}\neq 0\right\}$. By a similar argument to that in the proof of Theorem 3, under the condition on $\mu_{k}$ as stated in Theorem 4, we have $\widehat{\mathcal{S}}_{k}=\mathcal{S}_{k}$ for $k=1,\ldots,d$ with probability at least $1-\exp(C^{\prime}\log q)-\exp(C^{\prime\prime}\log pq)$. Note the diagonal element of the estimated projection matrix $\mathcal{P}(\widehat{\mathbf{B}})]_{jj}=0$ if and only if all one $\hat{\beta}_{jk}=0$ across $k=1,\ldots,d$. Hence, taking the union bound, we have $\mathcal{S}=\mathcal{S}$ with a probability at least $1-d\exp(C^{\prime}\log q)-d\exp(C^{\prime\prime}\log pq)$.
Appendix C Additional derivations
C.1 Proof of Lemma 3